### v2

# Utilizing Contextual Clues and Role Correlations for Enhancing Document-level Event Argument Extraction

Abstract
Document-level event argument extraction (EAE) is a vital but challenging task in the domain of information extraction. Most existing approaches focus on the interaction between arguments and event triggers, ignoring two critical points: the information of contextual clues and the semantic correlations among argument roles. In this paper, we propose the CARLG (Contextual Clues Aggregation and Role-based Latent Information Guidance) model, which consists of two modules: Contextual Clues Aggregation (CCA) and Role-based Latent Information Guidance (RLIG), effectively leveraging contextual clues and role correlations for improving document-level EAE. The CCA module adaptively captures and integrates contextual clues by utilizing context attention weights from a pre-trained encoder. The RLIG module captures semantic correlations through role-interactive encoding and provides valuable information guidance with latent role representation. Notably, our CCA and RLIG modules are compact, transplantable and efficient, which introduce no more than 1% new parameters and can be easily equipped on other span-base methods with significant performance boost. Extensive experiments on the RAMS, WikiEvents, and MLEE datasets demonstrate the superiority of the proposed CARLG model. It outperforms previous state-of-the-art approaches by 1.26 F1, 1.22 F1, and 1.98 F1, respectively, while reducing the inference time by 31%. Furthermore, we provide detailed experimental analyses based on the performance gains and illustrate the interpretability of our model.
keywords:
Event Argument Extraction (EAE), Information Extraction (IE), Natural Language Processing (NLP), Natural Language Understanding (NLU).
journal: Knowledge-Based Systems
1 Introduction
Within the domain of Information Extraction (IE), document-level event argument extraction (EAE) is a vital task focused on identifying event arguments and their corresponding roles within textual content associated with specific events. This task is a key process of event extraction (EE) xu2023multi ; lu2022explainable ; lv2022trigger and has broad implications across various fields. For instance, it is instrumental in the analysis of social media networks mitchell1974social ; DING2023110687 ; YANG2022109939 , enhancing recommendation systems  lietal2020gaia ; LI2024111124 ; DONG2022108954 , and improving dialogue systems  Zhang2020.08.03.20167569 ; ZHAO2023110927 . As illustrated in Fig. 1, with the event trigger “homicide” of event type Violence, the goal of EAE system is to identify argument “Freddie Gray” as the role victim, “Baltimore” as the place, and “Officer Caesar Goodson Jr.” as the killer.
Compared with sentence-level EAE liuetal2018jointly ; waddenetal2019entity ; tongetal2020improving , document-level EAE faces more complex issues, notably long-range dependencies and the need for inferences across multiple sentences, discussed in research ebner2020multi and lietal2021document . To solve these problems, some prompt-based methods ma2022prompt ; hsu2023ampere design suitable templates for long text, which solves the long-distance dependency problem and extract arguments in a generative slot-filling manner. Additionally, some works utilize graph structures employing heuristic approaches zhang2020two ; pouranbenveysehetal-2022-document , syntactical frameworks xuetal2021document ; liu2022document ; 9204800 , and Graph Neural Networks kipfsemi for the facilitation of reasoning across sentences. However, current mainstream approaches ignore two essential aspects: (a) the contextual clue information; (b) the semantic correlations among roles.
Contextual clues refer to text elements in the document, excluding the target arguments themselves, which provide beneficial guidance for predicting complex argument roles. These clues play a significant role in event argument extraction by offering valuable information. For instance, in Figure 1, the contextual clues, such as death, accused, and trial, serve as important indicators for identifying the arguments “Freddie Gray” and “Officer Caesar Goodson Jr” within the event Violence. By considering these contextual clues, the model can leverage wider context information to make more informed and accurate predictions. However, many prior studies xu2022two ; lietal2021document ; liu-etal-2023-enhancing-document just rely on the pre-trained transformer encoder to capture global context information in text implicitly, ignoring the necessity of context clues information that is highly relevant to the entity zhou2021document and target event ebner2020multi . Therefore, in this paper, we propose a Contextual Clues Aggregation (CCA) module, which leverages the contextual attention product from the pre-trained transformer encoder to aggregate the event-specific contextual clues for each argument, enhancing the candidate argument representation by incorporating broader and relevant contextual information.
Figure 1: A document from RAMS ebner2020multi dataset. The event violence has the trigger of Homicide and there are three distinct arguments, each assuming different roles. We use the color red to emphasize specific words, representing contextual clues that are significant for the extraction of arguments.
Within the context of real-world document-level EAE task, numerous roles exhibit semantic correlations. For example, as illustrated in Figure 1, there exists a strong semantic correlation between the roles of killer and victim, serving as beneficial guidance for extracting arguments associated with these roles within the target event Violence. In addition, many roles co-occur in multiple events, indicating strong potential for close semantic correlations among them. For a more in-depth analysis, we perform a frequency count and visualize the co-occurrence of the top 15 roles in the RAMS dataset, as shown in Figure 2. Notably, roles such as attacker, target and instrument show high co-occurrence frequency, suggesting their strong semantic correlations compared with other roles. Therefore, in this paper, we propose a Role-based Latent Information Guidance (RLIG) module, comprising role-interactive encoding, the fusion of role information and latent role guidance. Specifically, in role-interactive encoding, we add roles into the input sequence, enabling role embeddings to learn both semantic information and latent correlations among roles through self-attention mechanism. Subsequently, the latent role embeddings are combined with candidate argument spans via pooling and concatenation operations. Finally, we achieve alignment of argument spans, triggers, and roles within a unified space through a novel alignment mechanism utilizing Tucker Decomposition, tucker1964extension , which provides valuable information guidance for document-level EAE.
In summary, to better integrate and leverage our two techniques, we improve existing span-based baselines zhang2020two ; DBLP:journals/corr/abs-1904-05255 for document-level EAE, which has small parameter size and high efficiency. Then the CARLG (Contextual Clues Aggregation and Role-based Latent Information Guidance) model is proposed, by equipping CCA and RLIG modules on our baseline for the aforementioned two crucial points respectively. Significantly, each module leverages the fine-organized attentive weights obtained from the transformer-based pre-trained model, introducing less than 1% of new parameters. Additionally, these two modules can be easily applied to other span-based methods with transformer encoders and provide a significant performance boost in document-level EAE tasks (as shown in our Sec. 5.2), which are compact and transplantable. The computational cost of our model is also lower compared with the recent state-of-the-art approach TSAR, saving no less than 31% of inference time. ur contributions are summarized as follows:
  • 1.
    An enhanced span-based baseline for document-level EAE tasks is introduced, which demonstrates both parameter and computation efficiency, significantly outperforming existing span-based baselines
  • 2.
    We propose a Contextual Clues Aggregation (CCA) module, which adaptively identifies and aggregates the contextual clue information and enhances the representation of candidate arguments by incorporating event-specific contextual information.
  • 3.
    We introduce a Role-based Latent Information Guidance (RLIG) module designed to offer guidance based on latent role information, capturing the semantic correlations among roles.
  • 4.
    Extensive experiments on the public RAMS and WikiEvents lietal2021document datasets demonstrate that the proposed approach combined effectively with various span-based document-level EAE models and lead to significant performance improvements. Our CARLG consistently improves EAE performance on the biological dataset MLEE 10.1093/bioinformatics/bts407 , showing the generality of our approach in different domains.
Figure 2: The visualization represents the co-occurrence frequencies of the top 15 common roles within the RAMS test set, which is defined as the ratio between the co-occurrence times and the sum of the total occurrence times of two roles, and the self-co-occurrence count is set to zero.
2 Related Works
In this section, we provide a brief overview of the related works on the EAE task. We analyze the strengths and weaknesses of current mainstream EAE methods and emphasize the importance of considering contextual clues and role embedding in this task. Additionally, we discuss several applications in the NLP field that leverage contextual clues and relation embedding, which motivates our proposed approach.
2.1 Event Argument Extraction
Event argument extraction (EAE) is pivotal for improving event semantic understanding by identifying event arguments within a context, which constitutes a fundamental step in various downstream applications, including question answering liu2022document ; 6544211 ; 9257099 , text summarization allahyari2017text ; 8949697 , and assisted decision-making zhang2020effect ; 9339934 .
Previous approaches primarily concentrate on identifying the event trigger and its associated arguments from individual sentences. In chenetal2015event , a neural pipeline model was first proposed for event extraction, which was later extended by nguyengrishman2015event ; liuetal2017exploiting to incorporate recurrent neural networks (RNNs) and convolutional neural networks (CNNs) into the pipeline model.  liuetal2018jointly ; yanetal2019event utilize dependency trees to construct semantic and syntactic dependencies and relations, which captures the dependency of words in a sentence. waddenetal2019entity explores all possible spans and utilizes graph neural networks to construct span graphs for information propagation. Additionally, transformer-based pre-trained models  waddenetal2019entity ; tongetal2020improving ; liuetal2022dynamic achieve impressive performance in event extraction.
However, in real-world situations, elements of an event are often expressed across multiple sentences, leading to the emergence of document-level event extraction (DEE) 6910247 ; 7347400 ; SU2024111180 ; SCABORO2023110675 . DEE seeks to identify event arguments throughout whole documents, addressing the issue of extended dependencieswangetal2022query ; xu2022two . We summarize these recent DEE works into 3 categories: (1) Tagging-based methods. Works like wangetal2021cleve ; ducardie2020document apply the BiLSTM-CRF framework for document-level event extraction (DEE). Zheng zhengetal2019doc2edag introduces a mixed graph and a tracking module to capture the interrelations between events. (2) span-based methods. yangetal2021document proposes an encoder-decoder framework for parallel extraction of structured events. Works ebner2020multi ; zhang2020two predict argument roles for selected text segments under a maximum length restriction. Furthermore, Xu xu2022two suggests a dual-stream encoder complemented by an AMR-guided graph to manage extended dependencies. (3) generation-based methods. lietal2021document approaches the issue through conditional generation, whereas duetal2021grit views it as a sequence-to-sequence endeavor. Additionally, Wei weietal2021trigger redefines the task in terms of a reading comprehension challenge. Furthermore, Ren renetal2022clio integrates argument roles into document encoding to account for multiple role details for nested arguments. ma2022prompt ; liu2022document ; XU2023110375 propose a prompt-based method that utilizes slotted prompts for generative slot-filling to extract arguments.
Most existing document-level EAE methods focus on the interaction between arguments and event triggers, ignoring the correlations of roles among events, or the clue information provided by non-arguments context, which can be crucial for the task. Therefore, we attempt to enhance the document-level EAE model with useful contextual information and semantic role correlations.
2.2 Contextual Clues
Contextual clues refer to the information and cues present in the surrounding context of a given entity, event, or situation clark2016contextual ; zhang2020contextual ; liu2022document . These clues provide additional context and can help in understanding the meaning, significance, or relationships associated with the target entity or event peters2018deep . NC-DRE zhang2022nc introduces an encoder-decoder structure, leveraging a pre-trained encoding model and a graph neural network (GNN) for decoding. This setup facilitates the use of non-entity clues in document-level relation extraction. Concurrently, a method for localized context pooling zhou2021document has been put forward, allowing for the direct application of attention mechanisms from pre-trained language models. This method efficiently leverages the contextual information tied to particular entity pairs, aiding in capturing the beneficial context clues for determining relations. Xu xu2021discriminative uses the self-attention mechanism vaswani2017attention to learn a document-level context representation (DLCRep) which contributes to the improvement of information extraction.
Therefore, contextual clues contain important semantic information, which can provide information guidance for deep mining of semantic associations between entities. In this paper, we are the first to incorporate contextual clues techniques into the event argument extraction task and propose the Contextual Clues Aggregation (CCA) module to explicitly capture contextual clues.
2.3 Relation Embedding
Many methods in the field of information extraction leverage relation embeddings to enhance the effectiveness of various tasks 9086164 ; 9694521 ; 9123572 . CRE chen-badlani-2020-relation presents a novel approach to relation extraction by utilizing sentence representations as relation embeddings, which are assessed through entity embeddings trained concurrently with the knowledge base. xuetal2022emrel constructs relation representation explicitly and interacts it with entities and context, which achieves excellent results in multi-triple relation extraction.  han2022documentlevel employs relation embeddings as a medium and introduces two sub-tasks focused on co-occurrence prediction, aiming to enhance the retrieval of relational facts.
Relation embedding is typically utilized in relation extraction tasks. In the case of event argument extraction, we can consider event roles as relations connecting arguments and events. This viewpoint enables the utilization of relation embedding to proficiently capture the semantic correlations between arguments and events within EAE endeavors. Therefore, we aim to use contextual clues and role correlations embedded in latent embedding to strengthen the document-level EAE performance.
3 Enhanced Span-based Baseline
To better integrate our two techniques, we first propose an enhanced span-based baseline for document-level EAE, which exhibits small parameter size and high efficiency. We clarify the preliminaries of the document-level EAE task as follows for better understanding our method.
3.1 Preliminaries
We follow the definition of event argument extraction in RAMSv1.0 benchmark ebner2020multi . To better understand this task, we first clarify some important notions.
  • 1.
    Event mention: a text span within a document that mentions an entity object, e.g., the phrase Paul Duggan in Figure 1.
  • 2.
    Event trigger (trigger): the key term that most explicitly represents the manifestation of the event, e.g., the word homocide in Figure 1.
  • 3.
    Event type: the semantic class of an event, e.g., the event type Violence in Figure 1.
  • 4.
    Argument role (role): the relation between the event and the argument in which it is involved, for example, the role of Baltimore is identified as Place in Figure 1. For this task, argument roles are predetermined for every event type.
  • 5.
    Event argument (argument): an entity mention, value or temporal expression playing a specific argument role, e.g., the argument Freddie Gray in Figure 1.
Figure 3: The architecture or our enhanced baseline.
Formally, for a document $\mathcal{D}$ composed of $N$ words, denoted as $\mathcal{D}=\{w_{1},w_{2},\ldots,w_{N}\}$, a set of pre-defined event types $\mathcal{E}$, a corresponding set of argument roles $\mathcal{R}_{e}$ associated with each event $e\in\mathcal{E}$, and the event trigger $t$ in $\mathcal{D}$. The goal is to detect every pair $(r,s)$ for every event $e$ within the document $\mathcal{D}$, where $r\in\mathcal{R}_{e}$ signifies a role associated with the event $e$, and $s$ denotes a contiguous text span within $\mathcal{D}$. In alignment with earlier research xu2022two ; ebner2020multi , we extract event arguments independently for each event within a document.
3.2 Enhanced Span-based Baseline
In this section, we improve existing span-based baselines ebner2020multi ; zhang2020two and introduce a new baseline for document-level EAE. We integrate other techniques to further improve the performance and the architecture is shown in Figure 3.
3.2.1 Encoder
Following xu2022two ; ebner2020multi , we divide a document into several instances, where each instance only contains one event. For each document instance $\mathcal{D}$, the target event $e$ and the given trigger word $t_{e}$, we wrap the event trigger with special tokens "*""""*"" * " and regard the embedding of the start "*""""*"" * " token as the trigger embedding, which is adapted from the entity marker technique shi2019simple ; soares2019matching . We also add the event type as additional information guidance and the final input sequence $S$ is obtained as follows:
$S\ =\ \mathrm{[CLS]}\ [E_{e}]\ {e}\ [E_{e}]\ \mathrm{[SEP]}\ w_{1}\ w_{2}*\ t_{e}\ *\ ...\ w_{N}\ \mathrm{[SEP]},$
where $[E_{e}]$ is the special event token [^1] In our implementation, using [unused] tokens in BERT devlinetal2019bert and introducing specific tokens for RoBERTa DBLP:journals/corr/abs-1907-11692 . of event $e$. Next, we use a pre-trained transformer encoder to get the embedding of each token as follows:
$\mathbf{H}=\mathrm{Encoder}(\textsc{$S$}).$ (1)
Next, we can get the event representation $\mathbf{H}^{e}\in\mathbb{R}^{1\times d}$ for the start $[E_{e}]$ and the context representation $\mathbf{H}^{w}\in\mathbb{R}^{l_{w}\times d}$, where $l_{w}$ represents the length of the word pieces list. Notably, when the input sequence is longer than the maximum token length, we employ a dynamic window approach zhou2021document ; zhang2021document to process the long sequence. We divide the sequence into overlapping windows and calculate the embeddings for each window. Next, we obtain the final representation by averaging the embeddings of tokens that overlap across various windows.
3.2.2 Candidate Spans Representation
Filtering Impossible Spans   To eliminate noise information from useless spans, we implement a filtering step to reduce the number of candidate spans. Specifically, we exclude argument-impossible spans, such as spans containing commas in the middle. For a document $\mathcal{D}$, candidate spans after argument-impossible exclusion can be derived as $\mathcal{S}=\{s_{i:j}\}$. This enhancement allows us to eliminate approximately 25% of candidate spans on average, enabling our model to focus on candidate spans that contain valuable information.
Candidate Spans Representation  For a candidate argument span ranging from $w_{i}$ to $w_{j}$, we first obtain its average pooling representation $\textbf{s}_{i:j}$ as follows:
$\textbf{s}_{i:j}=\frac{1}{j-i+1}\sum_{k=i}^{j}\mathbf{h}_{k}^{w},$ (2)
where $\mathbf{h}_{k}^{w}$ is the $k^{\mathrm{th}}$ token embedding from $\mathbf{H}^{w}$.
As we extract arguments at the span level, where the boundaries can be ambiguous, we employ fully connected neural networks to construct start and end representations as follows:
$\mathbf{H}^{start}=\mathbf{W}^{start}\mathbf{H},$ (3)
$\mathbf{H}^{end}=\mathbf{W}^{end}\mathbf{H},$ (4)
where $\mathbf{H}$ represents the hidden representation of the input sequence $S$, and $\mathbf{W}^{start},\mathbf{W}^{end}\in\mathbb{R}^{d\times d}$ are trainable parameters. Subsequently, the candidate span representation $\widetilde{\mathbf{s}}_{i:j}$ can be obtained as follows:
$\widetilde{\mathbf{s}}_{i:j}=\mathbf{W}^{s}[\mathbf{h}^{start}_{i};\textbf{s}_{i:j};\mathbf{h}^{end}_{j}],$ (5)
where $\mathbf{W}^{s}\in\mathbb{R}^{3d\times d}$ is the trainable parameter and $\mathbf{h}^{start}_{i}$ and $\mathbf{h}^{end}_{j}$ are the $i^{\mathrm{th}}$ and $j^{\mathrm{th}}$ vector of $\mathbf{H}^{start}$ and $\mathbf{H}^{end}$. we establish the boundary loss for identifying the start and end positions in accordance with xu2022two :
$\begin{array}[]{c}\mathcal{L}_{b}=-\sum_{i=1}^{|\mathcal{D}|}[y_{i}^{s}\log P_{i}^{s}+(1-y_{i}^{s})\log(1-P_{i}^{s})\\ +y_{i}^{e}\log P_{i}^{e}+(1-y_{i}^{e})\log(1-P_{i}^{e})],\end{array}$ (6)
where $y_{i}^{s}$ and $y_{i}^{e}$ denote golden labels and $P_{i}^{s}=\mathrm{sigmoid}(\mathbf{W}_{1}\mathbf{h}^{start}_{i})$ and $P_{i}^{e}=\mathrm{sigmoid}(\mathbf{W}_{2}\mathbf{h}^{end}_{i})$ represent the probabilities of the word $w_{i}$ being predicted as the start or end word of a golden span.
3.2.3 Classification Module
Within each event $e$, for each candidate span $s_{i:j}$, we apply a comparison technique mou-etal-2016-natural ; zeng2020double to enhance its representation. In detail, we concatenate the span representation $\widetilde{\mathbf{s}}_{i:j}$ with the trigger representation $\mathbf{h}_{t}$, their absolute disparity $|\mathbf{h}_{t}-\widetilde{\mathbf{s}}_{i:j}|$, element-wise product $\mathbf{h}_{t}\odot\widetilde{\mathbf{s}}_{i:j}$, the embedding of the event type $\mathbf{H}^{e}$ and the embedding of the span length $\mathbf{E}_{len}$. Finally, through a feed-forward network, we calculate the predicted probability $P(r_{s_{i:j}})$ of the candidate span $s_{i:j}$:
$\mathbf{I}_{i:j}=[\widetilde{\mathbf{s}}_{i:j};\mathbf{h}_{t};|\mathbf{h}_{t}-\widetilde{\mathbf{s}}_{i:j}|;\mathbf{h}_{t}\odot\widetilde{\mathbf{s}}_{i:j};\mathbf{H}^{e};\mathbf{E}_{len}],$ (7)
$P(r_{s_{i:j}})=\mathrm{FFN}(\mathbf{I}_{i:j}).$ (8)
Considering the issue of imbalanced role distribution and the prevalence of negative samples among candidate arguments, we utilize focal loss 8237586 during the training process. Focal loss uses hyperparameters $\alpha$ and $\gamma$ to control the weighting and scaling of the loss function, which allows us to emphasize the importance of positive samples and makes the model focus more on useful information.
$\begin{array}[]{c}\mathcal{L}_{c}=-\sum_{s_{i:j}\in\mathcal{S}}\alpha{{[1-P(r_{s_{i:j}}=y_{s_{i:j}}}))]}^{\gamma}\\ \cdot\log P(r_{s_{i:j}}=y_{s_{i:j}}).\end{array}$ (9)
3.2.4 Training and Inference
During the training process, our model extracts single event at a time. Our training loss consists of two parts: classification loss $\mathcal{L}_{c}$ and boundary loss $\mathcal{L}_{b}$, and $\lambda$ is hyperparameter.
$\mathcal{L}=\mathcal{L}_{c}+\lambda\mathcal{L}_{b}.$ (10)
Consistent with the training process, we predict and choose the argmax role for each candidate span at inference time. If more than one argument span is predicted with the same role, we reserve all spans for this role at inference time.
Figure 4: The main structure of CARLG involves an input sequence, encompassing roles and event types, passing through a transformer-based encoder. This encoder generates output in the form of context representations, role representations, and attention heads. The Contextual Clue Aggregation (CCA) module dynamically consolidates contextual clues, while the Role-based Latent Information Generation (RLIG) module generates latent role embeddings and captures role correlations. Subsequently, the context vector is combined with the latent role vector to create the ultimate span representation. Finally, the classification module leverages this representation to forecast argument roles for all potential spans.
4 CARLG Model
In this section, we introduce Contextual Clues Aggregation (CCA) module which adaptively aggregates contextual clues and Role-based Latent Information Guidance (RLIG) module which effectively captures semantic correlations among roles. These two tecniques are then well equipped on our proposed baseline and our CARLG model are shown in Figure 4.
4.1 Contextual Clues Aggregation
The average pooling representation that previous span-based methods xu2022two ; ebner2020multi ; zhang2020two disregards important contextual information from non-argument words. Although the self-attention mechanism of the pre-trained encoder can capture token-level interactions, it lacks specific global interaction tailored to the event and candidate arguments. To address this limitation, we propose a Contextual Clues Aggregation (CCA) module that selects and combines relevant contextual information for each tuple comprising a candidate span and the event trigger word, denoted as $(s_{i:j},t)$. In this module, as shown in Figure 5, we leverage the attentive heads from the pre-trained transformer encoder, which enables the adoption of attentions from the pre-trained language model instead of training new attention layers.
We employ the token-level attention heads $\textbf{A}^{w}\in\mathbb{R}^{H\times l_{w}\times l_{w}}$ obtained from the final transformer layer of the pre-trained language model. Here, $H$ represents the number of attention heads, and $l_{w}$ signifies the length of word pieces. Utilizing these attention heads, we compute the context attention $\textbf{A}^{C}_{i:j}\in\mathbb{R}^{l_{w}}$ for each candidate span spanning from $w_{i}$ to $w_{j}$, by applying average pooling:
$\textbf{A}^{C}_{i:j}=\frac{1}{H(j-i+1)}\sum_{h=1}^{H}\sum_{m=i}^{j}\mathbf{A}^{w}_{h,m}.$ (11)
Then for span-trigger tuple $(s_{i:j},t)$, we get the contextual clue information $\textbf{c}^{s_{i:j}}\in\mathbb{R}^{d}$ which is crucial for the candidate span, by applying multiplication to the attentions and then normalizing the result:
Figure 5: Contextual Clues Aggregation Module.
$\begin{array}[]{c}\textbf{p}^{c}_{i:j}=softmax(\textbf{A}^{C}_{i:j}\cdot\textbf{A}_{t}^{C}\;)\vspace{1.3ex},\\ \textbf{c}^{s_{i:j}}={\mathbf{H}^{w}}^{T}\ \textbf{p}^{c}_{i:j},\end{array}$ (12)
where $\textbf{A}^{C}_{t}\in\mathbb{R}^{l_{w}}$ is the contextual attention of trigger $t$, $\cdot$ means dot product and $\textbf{p}^{c}_{i:j}\in\mathbb{R}^{l_{w}}$ is the computed contextual attention weight vector. $T$ is the transpose symbol.
We can apply our CCA module to other span-based methods by modifying Eq. 2 and fusing the contextual clue information $\textbf{c}^{s_{i:j}}$ into the average pooling representation as follows:
$\textbf{s}^{c}_{i:j}=tanh(\mathbf{W}^{CCA}[\frac{1}{j-i+1}\sum_{k=i}^{j}\mathbf{h}_{k}^{w};\textbf{c}^{s_{i:j}}]),$ (13)
where $\mathbf{W}^{CCA}\in\mathbb{R}^{2d\times d}$ is learnable parameter. $\textbf{s}^{c}_{i:j}$ is the enhanced representation of the candidate span $s_{i:j}$.
Figure 6: Role-based Latent Information Guidance Module. Sub-figure (a) illustrates the information fusion and (b) illustrates the Latent Role Information Guidance operation.
4.2 Role-based Latent Information Guidance
In this section, we propose Role-based Latent Information Guidance (RLIG) module, which is shown in Figure 6.
4.2.1 Role-interactive Encoding
To capture the semantic correlations between roles effectively, we integrate the information of role types into the input and utilize multi-head attention to facilitate interaction between the context and roles. This enables us to obtain the context and role representation in a shared knowledge space. We generate latent role embeddings by assigning unique special tokens   [^2] In our implementation, using [unused] tokens in BERT devlinetal2019bert and introducing specific tokens for RoBERTa DBLP:journals/corr/abs-1907-11692 to embody different roles in the framework of a pre-trained model, with each role type gaining a unique latent representation. Considering that role names carry valuable semantic information wangetal2022query , we encase role names in special tokens indicative of their role types, selecting the embedding of the initial special token as the role’s representation. This approach ensures the capture of unique attributes and subtle semantic distinctions of each role. For instance, the role “Victim” is denoted as $[R_{0}]\quad Victim\quad[R_{0}]$, with $[R_{0}]$ representing the specialized token for “Victim”.
Given the input document $\mathcal{D}$, the target event $e$ and its associated set of roles $\mathcal{R}_{e}=\{{r}_{1},{r}_{2},{r}_{3},...\}$, we enhance the input sequence described in Sec. 3.2.1 by concatenating these elements in the following manner:
$S\ =\ \mathrm{[CLS]}\ [E_{e}]\ {e}\ [E_{e}]\ \mathrm{[SEP]}\ w_{1}\ w_{2}\ *\ t_{e}\ *\ ...\ w_{N}\ \mathrm{[SEP]}$
$[R_{0}]\ {r_{0}}\ [R_{0}]\ [R_{1}]\ {r_{1}}\ [R_{1}]\ ...\ \mathrm{[SEP]}$,
where $[R_{0}]$ and $[R_{1}]$ denote the special role type tokens for $r_{0}$ and $r_{1}$, respectively, and we utilize the final $\mathrm{[SEP]}$ token to signify the ”none” category. By employing role-interactive encoding, we are able to get the event representation $\mathbf{H}^{e}\in\mathbb{R}^{1\times d}$ for the start event $[E_{e}]$, the context representation $\mathbf{H}^{w}\in\mathbb{R}^{l_{w}\times d}$, and the role representation $\mathbf{H}^{r}\in\mathbb{R}^{l_{r}\times d}$, all from $\mathbf{H}^{s}$. Here, $l_{w}$ denotes the length of the list of word pieces, and $l_{r}$ represents the length of the list of roles.
4.2.2 Role Information Fusion
Through role-interactive encoding, the role embeddings are able to capture semantic correlations and dynamically adapt to the target event and context. To ensure each candidate argument benefits from the guidance of role-specific information, we adapt our method for aggregating contextual clues to selectively incorporate role information. The aggregated role information $\textbf{r}^{s_{i:j}}\in\mathbb{R}^{d}$ for ${s_{i:j}}$ is obtained through contextual pooling, which involves adjusting the procedure described in Equations 11 and 12:
$\begin{array}[]{c}\textbf{A}^{R}_{i:j}=\frac{1}{H(j-i+1)}\sum_{h=1}^{H}\sum_{m=i}^{j}\mathbf{A}^{r}_{h,m}\vspace{1.3ex},\\ \textbf{p}^{r}_{i:j}=softmax(\textbf{A}^{R}_{i:j}\cdot\textbf{A}_{t}^{R}\;)\vspace{1.3ex},\\ \textbf{r}^{s_{i:j}}={\mathbf{H}^{r}}^{T}\ \textbf{p}^{r}_{i:j},\end{array}$ (14)
where $\textbf{A}^{r}\in\mathbb{R}^{H\times l_{w}\times l_{r}}$ represents the role attention heads obtained from the final transformer layer of the pre-trained model. $\textbf{A}^{R}_{i:j}\in\mathbb{R}^{l_{r}}$ represents the role attention for each candidate span, while $\textbf{A}^{R}_{t}\in\mathbb{R}^{l_{r}}$ denotes the role attention associated with the trigger $t$. Additionally, the vector $\textbf{p}^{r}_{i:j}\in\mathbb{R}^{l_{r}}$ characterizes the computed attention weight vector for roles. We can modify Eq. 13 and integrate the latent role information $\textbf{r}^{s_{i:j}}$ into the average pooling representation as follows:
$\textbf{s}^{r}_{i:j}=tanh(\mathbf{W}^{RLIG}[\frac{1}{j-i+1}\sum_{k=i}^{j}\mathbf{h}_{k}^{w};\textbf{r}^{s_{i:j}}]),$ (15)
where $\mathbf{W}^{RLIG}\in\mathbb{R}^{2d\times d}$ is learnable parameter.
4.2.3 Latent Role Information Guidance
We first map the original $d$-dimensional role features into a reduced dimension $d^{\prime}$ via a fully connected network to obtain the latent role information $\mathbf{H}^{r^{\prime}}\in\mathbb{R}^{l_{r}\times d^{{}^{\prime}}}$ as follows:
$\mathbf{H}^{r^{\prime}}=\mathbf{H}^{r}\mathbf{W}_{3}+\mathbf{b}_{w},$ (16)
where $\mathbf{W}_{3}\in\mathbb{R}^{d\times d^{{}^{\prime}}}$ and $\mathbf{b}_{w}\in\mathbb{R}^{d^{{}^{\prime}}}$ are learnable parameters. To fully utilize the expressive power of role embeddings, we propose a factorization-based alignment approach using Tucker decomposition  tucker1964extension . We define candidate span-trigger-role triples set $\mathcal{T}=\{<s_{i:j},r_{k},t_{e}>|s_{i:j}\in\mathcal{S},r_{k}\in\mathcal{R}_{e}\}$. Then we introduce a core tensor $\mathbf{Z}\in\mathbb{R}^{d_{i}\times d^{{}^{\prime}}}$ where $d_{i}$ is the hidden dimension of the combined representation $\mathbf{I}_{i:j}$ calculated in Eq. 7, and for simplity, the validity for each triple $<s_{i:j},r_{k},t_{e}>\in\mathcal{T}$ is scored as follows:
$\varphi(s_{i:j},r_{k},t_{e})\;=\;\sigma(\mathbf{Z}\;\times_{1}\;\mathbf{I}_{i:j}\times_{2}\;{\widehat{\mathbf{r}}}_{\mathrm{k}}\;+b_{k}),$ (17)
where ${\widehat{\mathbf{r}}}_{\mathrm{k}}=\mathbf{H}^{r^{\prime}}_{k,:}$ represents the transformed role features and ×$\times_{n}$ denotes the tensor product operation along the $n^{th}$ dimension. The calculation of $\varphi$ for all triples is performed concurrently via a batched tensor product technique, and the model is trained employing focal loss as described in reference 8237586 .
5 Experiments
5.1 Experimental Setup
5.1.1 Datasets
We evaluate the proposed method on three large-scale document-level EAE datasets, that is, RAMS ebner2020multi , WikiEvents lietal2021document and MLEE 10.1093/bioinformatics/bts407 , and detailed data statistics are listed in Table 1.
Table 1: Detailed data statistics.
Dataset RAMS WikiEvents MLEE
# Event types 139 50 23
# Args per event 2.33 1.40 1.29
# Events per text 1.25 1.78 3.32
# Role types 65 57 8
# Events
 Train 7329 3241 4442
 Dev 924 345 -
 Test 871 365 2200
  • 1.
    RAMS  The dataset includes 9,124 annotated events sourced from English online news articles. In this dataset, each event forms a single instance, and events that occur within the same context are grouped together into a single instance containing multiple events. We adopt the original train/dev/test split following ebner2020multi .
  • 2.
    WikiEvents  This dataset centers on events documented in English Wikipedia, encompassing both the events and the related news articles referencing these events. It’s annotated with co-reference links for arguments. However, for our experiments, we only employ the precise argument annotations provided.
  • 3.
    MLEE  This dataset is a document-level event extraction dataset that focuses on bio-medical publications, consisting of manually annotated abstracts written in English. We follow the procedure outlined in 10.1093/bioinformatics/btaa540 and it’s worth noting that the preprocessed dataset only provides a train/test data split, without a dedicated development set. In this case, we use the training set as the development set for our experiments following  10.1093/bioinformatics/btaa540 .
5.1.2 Evaluation Metrics
For the RAMS dataset, we adopt the method detailed in xu2022two , focusing on both Span F1 and Head F1 scores across the development and test sets. Span F1 evaluates the precision of matching predicted argument spans against the gold standard, while Head F1 quantifies the accuracy of identifying the central word in each argument span. For a comprehensive assessment, we also incorporate Span F1 for argument identification (Arg IF) on the test set, as suggested by ma-etal-2022-prompt . Regarding the the WikiEvents dataset, we present both Head F1 and Coref F1 metrics on the test set for Argument Identification (Arg IF) and Argument Classification (Arg CF) tasks, consistent with the methodology described in lietal2021document . The Coref F1 score, inspired by the evaluation technique from jigrishman2008refining , gauges the coreference accuracy between extracted arguments and the established gold standard. For the MLEE dataset, we report Span F1 scores on the test set for both Arg IF and Arg CF tasks, following the methodologies in lietal2021document ; ma-etal-2022-prompt .
5.1.3 Baselines
[…]
Table 2: Detailed hyperparameter settings.
Hyperparameters RAMS WikiEvents MLEE
Training Epochs 50 100 100
Warmup Ratio 0.2 0.05 0.1
Transformer LR 3e-5 3e-5 3e-5
Not Transformer LR 1e-4 1e-4 1e-4
Dropout Ratio 0.1 0.1 0.1
Batch Size 4 4 4
Max Span Length 8 8 8
Max Input Length 1024 1024 1024
Lambda 0.05 0.1 0.05
Alpha 10 10 10
Gamma 2 2 2
5.1.4 Hyperparameter Setting
Due to the fact that our proposed module does not rely heavily on hyperparameters, most of the hyperparameters follow the same configuration of TSAR xu2022two . More detailed hyperparameters settings are listed in Table 2.
Table 3: The main results for the RAMS dataset. Arg IF denotes the F1 score for the Argument Identification subtask, and other metrics are for the Argument Classification subtask. * indicates results reported by  ma-etal-2022-prompt . The highest scores are highlighted in bold, while the second-highest scores are underlined.
Method Dev Test
Span F1 Head F1 Span F1 Head F1 Arg IF
Tag-based
BERT-CRF ebner2020multi 38.1 45.7 39.3 47.1 -
[…]
Generation-based
FEAE weietal2021trigger - - 47.40 - -
[…]
EACE zhou2024eace - - 50.1 56.3
RKDE hu2023role - - 50.3 55.1
Span-based
Two-Step zhang2020two 38.9 46.4 40.1 47.7 -
[…]
${\text{TSAR}}_{{\text{base}}}$ xu2022two 45.23 51.70 48.06 55.04 -
${\text{TSAR}}_{{\text{large}}}$ xu2022two 49.23 56.76 51.18 58.53 -
${\text{SCPRG}}_{{\text{base}}}$ liu-etal-2023-enhancing-document 46.06 52.75 48.18 55.83 -
${\text{SCPRG}}_{{\text{large}}}$ liu-etal-2023-enhancing-document 50.23 57.22 51.88 59.06 -
Our Methods
[…]
${\text{CARLG}}_{{\text{base}}}$ 46.83 53.42 49.81 56.21 55.91
${\text{CARLG}}_{{\text{large}}}$ 50.64 57.71 52.82 59.79 56.71
Span-based+Ours
Two-Step+Ours 43.42 50.58 45.27 52.79 50.34
${\text{TSAR}}_{{\text{base}}}$+Ours 46.52 53.28 49.23 56.07 53.22
${\text{TSAR}}_{{\text{large}}}$+Ours 50.61 57.96 52.52 59.77 57.03
5.2 Main Results
5.2.1 Performance of our enhanced baseline
[…]
5.2.2 Performance of CARLG
[…]
Table 4: Main results of WikiEvents and MLEE datasets. * means we rerun their code based on their experimental settings (hyperparameter tuning is conducted when necessary). The highest scores are highlighted in bold, while the second-highest scores are underlined.
WikiEvents MLEE
Method Arg IF Arg CF Arg IF Arg CF
Head F1 Coref F1 Head F1 Coref F1 Span F1 Span F1
Tag-based
BERT-CRF ebner2020multi 69.83 72.24 54.48 56.72 - -
Generation-based
BERT-QA du-cardie-2020-event 61.05 64.59 56.16 59.36 - -
BERT-QA-Doc du-cardie-2020-event 39.15 51.25 34.77 45.96 - -
EEQA du-cardie-2020-event - - 56.9 - 68.4 66.7
[…]
${\text{EA}^{2}\text{E}}_{{\text{large}}}$ zengetal2022ea2e 74.62 75.77 68.61 69.70 - -
EACE zhou2024eace - 71.1 - 66.2 - -
RKDE hu2023role - 69.1 - 63.8 - -
Span-based
Two-Step zhang2020two 70.32* 72.27* 55.11* 57.12* 70.64* 68.74*
${\text{TSAR}}_{{\text{base}}}$ xu2022two 75.52 73.17 68.11 66.31 72.21* 71.03*
${\text{TSAR}}_{{\text{large}}}$ xu2022two 76.62 75.52 69.70 68.79 73.92* 72.71*
${\text{SCPRG}}_{{\text{large}}}$ liu-etal-2023-enhancing-document 77.52 76.01 70.11 69.25 73.88* 73.21*
Our Methods
[…]
${\text{CARLG}}_{{\text{base}}}$ 76.22 74.97 69.03 68.53 73.26 71.23
${\text{CARLG}}_{{\text{large}}}$ 77.22 76.03 70.92 70.11 75.90 73.71
Span-based+Ours
Two-Step+Ours 73.20 72.96 60.90 61.04 72.42 70.11
${\text{TSAR}}_{{\text{base}}}$+Ours 76.43 75.10 68.81 68.22 73.13 71.12
${\text{TSAR}}_{{\text{large}}}$+Ours 77.46 76.10 70.88 70.28 74.98 73.48
5.2.3 Performance of Span-based Methods equipped with CARLG
In order to evaluate the compatibility of our CCA and RLIG modules, we equip these two modules on some classical span-based methods and conduct experiments on three datasets. As is shown in Table 3 and Table 4, both Two-Step and TSAR exhibit improved performance when equipped with CCA and RLIG modules. Two-Step yields an improvement of +5.17/+5.09 Span F1/Head F1 for Arg CF on RAMS, +2.88/+5.79 Head F1 for Arg IF/Arg CF on WikiEvents, and +1.78/+1.37 Span F1 for Arg IF/Arg CF on MLEE. ${\text{TSAR}}_{{\text{large}}}$ achieves notable improvements in Span F1/Head F1 for Arg CF on RAMS (+1.34/+1.24), Head F1 for Arg IF/Arg CF on WikiEvents (+0.84/+1.18), and Span F1 for Arg IF/Arg CF on MLEE (+1.06/+0.77). It is worth noting that the improvement of TSAR after incorporating CCA and RLIG modules is relatively modest. This could be attributed to the fact that the AMR graph already contains contextual semantic information, which may mitigate the need for additional contextual clue integration. These results illustrate that our two modules can be easily applied to other span-based methods with transformer encoders and provide a significant performance boost in document-level EAE tasks, which are compact and transplantable.
[…]
Method Dev Test
Span F1 Head F1 Span F1 Head F1
[…]
 -Boundary Loss 42.18 50.28 42.33 51.68
 -Focal Loss 44.97 51.36 45.83 53.03
 -FIS 43.27 50.64 45.26 52.11
Table 6: Ablation Study on RAMS and WikiEvents for CARLG. Span, Head and Coref refer to Span F1, Head F1 and Coref F1 scores respectively.
RAMS WikiEvents
Method Params Dev Test Arg IF Arg CF
Span Head Span Head Head Coref Head Coref
${\text{CARLG}}_{{\text{base}}}$ 120.87M 46.83 53.42 49.81 56.21 76.22 74.97 69.03 68.53
-CCA 120.28M 45.78 52.67 47.43 54.57 74.57 73.86 67.44 66.87
-RLIG 118.78M 45.27 52.36 47.91 54.52 75.59 73.83 68.26 67.49
-both 118.19M 45.07 51.59 45.76 53.16 73.92 73.22 66.90 65.98
${\text{CARLG}}_{{\text{large}}}$ 374.70M 50.64 57.71 52.82 59.71 77.22 76.03 70.92 70.11
-CCA 373.65M 49.68 56.46 50.84 57.9 75.68 73.82 69.79 68.65
 -RLIG 371.68M 49.96 57.32 51.28 58.46 76.30 74.08 69.87 68.96
-both 370.63M 48.33 54.04 47.52 55.61 75.45 73.55 68.63 67.30
5.3 Ablation Study
In order to effectively showcase the effectiveness of our components, we carry out an ablation study on CARLG using the RAMS and WikiEvents datasets, as depicted in Table 6. First, when we remove the Contextual Clues Aggregation (CCA) module, we observe a decrease in Span F1 and Head F1 scores of ${\text{CARLG}}_{{\text{base}}}$/ ${\text{CARLG}}_{{\text{large}}}$ by 2.83/1.98 and 1.64/1.81 on the test set, respectively, which indicates that our CCA module plays a crucial role in capturing essential contextual clues for document-level EAE.
Moreover, when removing the Role-based Latent Information Guidance (RLIG) module [^3] We additionally remove the respective role tokens that were included in the input sequence., we observe a significant decrease in performance for ${\text{CARLG}}_{{\text{base}}}$/ ${\text{CARLG}}_{{\text{large}}}$, with a drop of 1.9/1.54 in Span F1 and 1.69/1.07 in Head F1 on the RAMS test set. This demonstrates the effect of the RLIG module in capturing meaningful semantic correlations among roles, improving the performance of document-level EAE. Moreover, when both the CCA and RLIG modules are removed, the decline in performance is even more noticeable in comparison to excluding a single module. This illustrates the synergistic enhancement achieved by the collaboration of these two modules.
[…]
Figure 7: Context weights for a sample from RAMS are displayed, focusing on the span-trigger pair (Islamic State, wounded). Various color intensities are used to depict the attention weights.
5.4 Analysis of Contextual Attentive Weights
To evaluate the validity of our Contextual Clues Aggregation (CCA) module in capturing beneficial contextual clues for candidate argument spans, we present a visualization of the contextual weights $\textbf{p}^{c}_{i:j}$, as detailed in Eq.12, using an instance from the RAMS dataset. As shown in Figure 7, the CCA module assigns high weights to non-argument contextual clues such as attack, responsibility, and terrorist attack, which are highly relevant to the span-trigger pair (Islamic State, wounded). Notably, our Contextual Clue Aggregation (CCA) module also grants relatively high attentive weights to some words in other arguments, including explosive, Dozens, and Kabul. This implies that these argument words offer crucial information for predicting the role of Islamic State. This visualization illustrates how our Contextual Clue Aggregation (CCA) module effectively identifies relevant contextual clues for candidate spans and models the interaction of information among interconnected arguments within an event.
Figure 8: We provide visualizations of attention weights directed towards the context for various candidate spans within an event.
Additionally, our investigation extends to examining the attention weights associated with various span-trigger pairings within a single event. In Figure 8, we display a heat map created from the attention weights of 30 randomly chosen candidate spans in an event, with respect to their contextual relevance. The heat map reveals that different candidate arguments pay attention to different contextual information, suggesting that our Contextual Clues Aggregation (CCA) module is capable of dynamically identifying and utilizing pertinent contextual clues specific to each candidate argument span. This adaptability highlights the module’s efficiency in tailoring its focus to suit the unique contextual requirements of each argument span within an event.
Figure 9: The visualization of cosine similarity between role representations of two examples from RAMS in sub-figure (a) and MLEE dataset in sub-figure(b) respectively.
Figure 10: Here’s an instance of a t-SNE visualization sourced from the RAMS dataset. The representations of roles and arguments originate from five distinct documents. In sub-figure (a), we utilize average pooling representations encoded by BERT, while in (b), we visualize the representations fused with latent role embeddings.
5.5 Analysis of Role Correlations
To verify the capability of our model in capturing semantic correlations between roles, we visualize the cosine similarity among latent role representations from events within the RAMS and MLEE datasets. Figure 9 presents the results, where roles such as origin and destination, attacker and target, ToLoc and FromLoc exhibit similar representations, indicating their semantic similarity. This visualization validates our model’s efficacy in capturing semantic correlations among roles.
Additionally, to substantiate the effective guidance provided by role representations, we utilize t-SNE visualization JMLR:v9:vandermaaten08a to display arguments corresponding to two distinct roles that appear together in five diverse documents, alongside their respective latent role embeddings. Figure 10a illustrates that arguments associated with the same role but in distinct documents are dispersed throughout the embedding space. This dispersion is attributed to variations in the target events and their contextual settings. However, after integrating latent role embeddings, the representations of arguments linked to identical roles, such as victim or place, cluster more closely. This shows that the Role-based Latent Information Guidance (RLIG) module effectively provides crucial role-specific information. The visualization demonstrates that the RLIG module enhances the alignment of arguments within the same role across different contexts, thereby affirming its utility in providing insightful guidance of roles.
5.6 Complexity and Efficiency
Our CARLG method combines computational efficiency with low complexity while both CCA and RLIG introduce few parameters. As shown in Table 6, for ${\text{CARLG}}_{{\text{large}}}$, CCA utilizes the pre-trained encoder’s attention heads and performs multiplication and normalization operations, adding only about 0.28% new parameters. RLIG introduces approximately 0.81% new parameters totally. This makes our parameter quantity approximately equivalent to the transformer-based encoder plus a MLP classification layer.
Table 7 reports the efficiency comparison of various span-based models on MLEE dataset. Our ${\text{CARLG}}_{{\text{base}}}$ is approximately 2 times faster than ${\text{TSAR}}_{{\text{base}}}$ and can save up to 30.9% of inference time with smaller model size. Additionally, compared to Two-Step, our CARLG only increases 15.9% inference time and 8% model parameters. Such high efficiency and performance makes CARLG a promising choice for practical applications of document-level EAE.
Table 7: Time cost (in seconds) of different span-based models on MLEE dataset. Experiments are run for one epoch on one Tesla A100 GPU.
Model Params Training Time Inference Time
Two-Step 111.96M 229.68 55.29
${\text{TSAR}}_{{\text{base}}}$ 142.03M 465.84 92.79
${\text{CARLG}}_{{\text{base}}}$ 120.87M 280.08 64.08
5.7 Error Analysis
Table 8: Error analysis on the RAMS test set. The best results are in bold font.
Model Wrong Span Over-extract Partial Overlap Wrong Role Total
Two-Step 86 64 47 32 46 275
${\text{TSAR}}_{{\text{base}}}$ 81 48 57 28 19 233
${\text{CARLG}}_{{\text{base}}}$ 79 45 55 29 15 223
Figure 11: An extraction case from RAMS test set base on ${\text{Roberta}}_{{\text{large}}}$. This case shows both CCA module and RLIG module contribute to the argument extraction.
To conduct a more comprehensive examination of the errors made by various models and analyze the underlying reasons for these errors , we manually compared the predictions of 200 randomly selected examples from the RAMS test set against their golden annotations. During our analysis, we categorized the errors into five main categories following xu2022two . Wrong Span refers to the case where a specific role is assigned to a span that does not overlap with the corresponding golden span. We observe that this type of errors often occur when negative words such as ”not” are involved or when there are coreference spans related to the golden span. The category labeled Over-extract pertains to instances where the model predicts an argument role that does not genuinely appear in the document. Some of the extracted spans are either sub-strings of the golden spans (Partial) or exhibit overlaps with them (Overlap). The ”Partial” error category is also commonly observed when there is punctuation, such as a comma, within the golden span. Wrong Role refers to the model correctly identifies the golden span, but assigns the wrong argument role to it. We compare the errors of some span-based methods, that is, Two-Step and ${\text{TSAR}}_{{\text{base}}}$ and ${\text{CARLG}}_{{\text{base}}}$. As shown in Table 8, compared with Two-Step and ${\text{TSAR}}_{{\text{base}}}$, our CARLG model decreases the number of errors from 275 to 223 and effectively reduces Wrong Span errors, which can be attributed to the integration of contextual clues and latent role guidance.
Method RAMS
Arg-I Arg-C
text-davinci-003 46.1 39.5
gpt-3.5-turbo 38.3 31.5
gpt-4 50.4 42.8
CARLG (Ours) 56.71 52.82
Table 9: Comparison with large language model methods on the test set of RAMS.
5.8 Comparison with Large Language Models
Large language models (LLMs) have become a focal point of interest among researchers, given their remarkable versatility across a diverse range of Information Extraction (IE) tasks xu2023large ; li2023semi ; zhang2024ultra . In this paper, we undertake a comprehensive comparison with some large language models on the document-level EAE task. Due to the high cost associated with using GPT-4, its evaluation is limited solely to the RAMS dataset. The experimental results are detailed in Table 9. We utilize the prompts defined in the work zhou2023heuristics and experiments are conducted utilizing three prominent large language models: text-davinci-003 ouyang2022training , gpt-3.5-turbo ouyang2022training and GPT-4 achiam2023gpt , all of which are accessed via OpenAI’s public APIs  [^4] https://openai.com/api/. As illustrated in Table 9, LLMs exhibit a notable performance disparity compared to supervised learning models in the EAE task. Furthermore, the operational costs associated with large models are inherently substantial. In contrast, our approach demonstrates superior efficiency, cost-effectiveness, and overall effectiveness in the document-level EAE task when compared to LLMs.
5.9 Case Study
In order to explore how the CCA and RLIG modules work in actual cases, we conduct the case study for our CARLG. As shown in Figure 11, equipped with CCA and RLIG modules, our CARLG can correctly predict all arguments, while TSAR fails to predict arguments attack. When removing the CCA module, our CARLG fails to predict the argument Talha Hesham Abdo, indicting the CCA module contributes to the argument extraction by providing essential contextual clues. Additionally, when the RLIG module is removed, it fails to predict the argument for the crim role and mispredicts the argument for the jailer role. This can be attributed to that the RLIG module effectively captures the semantic correlations among the roles jailer, detainee, and crim and provides valuable information guidance, allowing the model to make more accurate predictions.
6 Conclusion
In this paper, we introduce the CARLG model for document-level event argument extraction (EAE). CARLG comprises two novel modules: Contextual Clues Aggregation (CCA) and Role-Based Latent Information Guidance (RLIG). CCA effectively aggregates non-argument clue words to capture beneficial contextual information, while RLIG provides guidance based on latent role representations that capture semantic correlations between some roles. Diverse experiments show that our CARLG model surpasses current SOTA EAE models on three large-scale document-level datasets.
While our CARLG model has exhibited remarkable performance in document-level event argument extraction (EAE) tasks with known event triggers, we acknowledge its limitation in depending on the availability of event triggers. In practical situations, event triggers may not always be supplied or known in advance. Therefore, in our future research, we intent to transfer our approach to encompass other EAE tasks that do not rely on explicit trigger words. We aim to conduct extensive experiments to assess its performance in such scenarios. Furthermore, we intend to explore the applicability of the CARLG framework in other information extraction tasks, such as relation extraction.